The displacement vectors A and B, when added together, give the resultant vector R, so that R = A + B. Use the data in the drawing and the fact that φ = 28° to find the magnitude R of the resultant vector and the angle θ that it makes with the +x axis.
1
Expert's answer
2019-10-07T10:41:33-0400
"R=\sqrt{6^2+8^2+2(6)(8)\cos{(90+28)\degree}}=7.4"

We have:


"8\cos{28\degree}=7.4\cos{\theta}"

"\theta=\arccos{0.9545}=17\degree"